• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name jpchen

    Log entry time 00:10:10 on April 4,2008

    Entry number 229825

    keyword=Swing shift summary

    Swing shift summary:

    Shift workers: J. P. Chen (SL), J. Huang (TO).

    Shift started with no beam.
    16:20 beam back. requested 5 uA, raster off.
    BPM poosition fine. requested harp scan.
    16:35, beam off.
    16:45 beam back.
    16:50 harp scan completed
    16:56 moved to BeO.
    17:00 asked for raster on 1x1 (MCC 1.5x 1)
    17:12, moved to LH2, asked for 3 uA.
    17:16 elastic run on LH2 to check detectors/tracking, left run 2257.
    17:18 right coda could not start.
    17:20 beam off, tuning for Hall B. left coda also crashed. BigBIte HV tripped
    17:25 restarted left daq.
    17:26, re-stared right daq.
    17:30 Disable and re-enable BigBite HV, back to 1400V for all,
    17:35 beam back.
    17:36 ROC5 crashed again. Restarted left coda. Start data taking.
    HV=1400 V, Threshold= 5V. 2259-2262
    17:54 Scintillator HV ch L1.4 tripped, reset.
    17:57 a few channles of Scientillators tripped: L1.4, L5.7, L7.9 then L7.1. Reset.
    18:05 raised threshold from 5 V to 7.5 V for both front and back chambers, run 2263
    18:15 DAQ crashed.
    requested raster off
    18:20 run 2264
    18:28 raised threshold to 10V. run 2265
    18:41 raised threshold to 12.5V, run 2266
    19:00 MCC took beam away to tune to get rid of scraping.
    122:08 lowered threshold to 7.5V, run 2279:00 change BigBIte Chamber HV from 1400 V to 1350V.
    21:10 beam back, start run 2270.
    22:00 lowered threshold to 10 V, run 2271.
    22:08 lowered threshold to 7.5V, run 2272.
    22:14 lowered threshold to 5V, run 2273.
    22:20 MCC took beam away to investigate problem in the injector.
    22:22 changed HVs from 1350 to 1450 V.
    23:45 beam back. run 2274.
    23:54 raised threshold from 5 to 7.5 V. run 2275.
    24:01 raised threshold to 10 V, run 2276.
    24:0 raised threshold to 12.5. run 2277, completion of scan.

    run 2257, first run, testing.
    Runs for threshodl/HV scans, most runs are 0.5M events:
    run 2259/2262, threshold=5V, HV=1400 V.
    run 2263, threshold=7.5V, HV=1400V, (DAQ crashed)
    run 2264, threshold=7.5V, HV=1400V, raster off.
    run 2265, threshold=10V, HV=1400V.
    run 2266, threshold=12.5V, HV=1400V.
    run 2270, threshold=12.5V, HV=1350V. 4.4M events.
    run 2271, threshold=10V, HV=1350V.
    run 2272, threshold=7.5V, HV=1350V.
    run 2273, threshold=5V, HV=1350V.
    run 2274, threshold=5V, HV=1450V.
    run 2275, threshold=7.5V, HV=1450V.
    run 2276, threshold=10V, HV=1450V.
    run 2277, threshold=12.5V, HV=1450V.